# Extracting Resonance Width from Lattice Quantum Monte Carlo Simulations Using Analytical Continuation Method

### 26 Mar 2026 v1

@@ 1 Introduction
Nuclear Lattice Effective Field Theory (NLEFT) is an efficient ab initio method utilizing the auxiliary-field quantum Monte Carlo (QMC) approach. Significant progress has been made with NLEFT in recent years, enabling studies of nuclear ground and excited states EPJA31-105 , PRL104-142501 , EPJA45-335 , PLB732-110 , PRL112-102501 , PLB869-139839 , PRL134-162503 , clustering phenomena PRL106-192501 , PRL109-252501 , PRL110-112502 , PRL119-222505 , hypernuclei EPJA60-215 , PRD111-036002 , scattering Nature528-111 , PRC86-034003 , JPGPP52-045103 , thermodynamics PRL115-185301 , and electro-weak reactions PLB859-139086 , PRC112-025502 . Within NLEFT, the nuclear Hamiltonian is discretized on a spatial cubic lattice and solved using imaginary-time projection to obtain low-lying eigenstates. Most NLEFT studies to date have focused on the bound states or treated very narrow resonances as the bound states approximately. For example, the energy of the second $0^{+}$ state in 12C (Hoyle state), which plays a crucial role in stellar nucleosynthesis Hoyle1954 , lies just above the three-$\alpha$ threshold. It is a resonance with an extremely narrow width of a few electron-volts against decay into three $\alpha$-particles. In such cases, direct NLEFT simulations can reliably extract its energy PRL106-192501 , PRL109-252501 , PRL110-112502 . However, these calculations provide no information on the resonant width and cannot be directly applied to broad resonances. So far, precisely describing general resonance properties remains particularly challenging within NLEFT.
@@
Generally, the single-particle resonances in the potential model correspond to the poles of $S$-matrix. A variety of theoretical approaches have been established to extract resonance parameters, $\it{i.e.}$ resonance energy $E$ and width $\Gamma$. These include the $R$-matrix and $K$-matrix formalisms PR72-29 , PRL59-763 , PRC44-2530 , Taylor2012 , PRC66-024311 , PRL109-072501 , PRC88-024323 , PRC81-034311 based on fitting to scattering or reaction observables, the Green’s function techniques SJNP45-783 , PRC90-054321 , CPC44-084105 , the complex scaling method (CSM) PRC34-95 , PRC37-383 , PRL79-2217 , the complex momentum representation method PRC73-034321 , PRL117-062502 , PLB801-135174 , NST2023Xu , the real stabilization method PRA1-1109 , PRC77-014312 and the phase-shift method PRC2003Sandulescu suitable for very narrow resonance, as well as the practical analytical continuation in the coupling constant (ACCC) method especially suitable for broad resonances PRC70-034308 , EPJA48-40 , PLB730-30 , PRC92-024324 , CCP21-1154 , PLB850-138511 , CPC49-014108 , PRC106-064320 , PRC113-014316 , etc.
For few‑body systems, some theoretical methods have been developed to describe resonance states from realistic or chiral nuclear interactions. The complex scaling method and the Jost function approach map resonances into isolated poles in the complex energy plane and enable the calculations of unbound states without explicit continuum discretization hale1987 , rakityansky2022 . The Gamow shell model has been used to study the resonances starting from the bare nuclear forces PRC87-034326 , PRC88-044318 , PRC99-061302 , PRC104-024319 . For ab initio few‑body calculations, the Faddeev–Yakubovsky (FY) formalism has achieved high-precision predictions for resonances in light nuclei such as 5He using modern chiral two- and three-nucleon forces lazauskas2019 , deltuva2007 . An ab initio approach to open quantum systems has been implemented for light nuclei by using a Gamow–Hartree–Fock basis and realistic interactions and employing coupled-cluster theory to solve the quantum many-body problem for the helium chain PLB656-169 . Other ab initio methods like the Green’s function Monte Carlo (GFMC) Carlson2005 or the no-core shell model PRC-Caurier2006 have been used to calculate the structure of helium isotopes, which allows nonperturbative calculations with realistic potentials but often shows larger deviations for resonance observables  wiringa1995 .
@@
In lattice quantum chromodynamics and finite‑volume ab initio calculations, the Lüscher method relates scattering phases and resonance parameters to energy levels in a finite box, providing a first‑principles constraint on few‑body resonances lattice2025 . Resonances appear as plateaus, with slopes indicating widths. This finite-volume approach was recently applied in NLEFT using Wigner-SU(4) and N3LO chiral forces to study the plausible tetraneutron resonance arXiv:2601.01801 . An efficient alternative involves computing scattering-amplitude poles in the complex-energy plane, where pole positions directly yield resonance energies (real parts) and widths (imaginary parts). These methods provide more direct resonance signatures and better resolve complex structures like overlapping poles or broad resonances. However, they require analytical continuation from real to imaginary momentum, which is a typically ill-posed problem introducing significant numerical uncertainties. Precise calculations thus demand both high-accuracy numerics and efficient analytical-continuation algorithms.

### 01 Apr 2026 v2

@@ 1 Introduction
Nuclear Lattice Effective Field Theory (NLEFT) is an efficient ab initio method utilizing the auxiliary-field quantum Monte Carlo (QMC) approach. Significant progress has been made with NLEFT in recent years, enabling studies of nuclear ground and excited states EPJA31-105 , PRL104-142501 , EPJA45-335 , PLB732-110 , PRL112-102501 , PLB869-139839 , PRL134-162503 , clustering phenomena PRL106-192501 , PRL109-252501 , PRL110-112502 , PRL119-222505 , hypernuclei EPJA60-215 , PRD111-036002 , scattering Nature528-111 , PRC86-034003 , JPGPP52-125102 , thermodynamics PRL115-185301 , and electro-weak reactions PLB859-139086 , PRC112-025502 . Within NLEFT, the nuclear Hamiltonian is discretized on a spatial cubic lattice and solved using imaginary-time projection to obtain low-lying eigenstates. Most NLEFT studies to date have focused on the bound states or treated very narrow resonances as the bound states approximately. For example, the energy of the second $0^{+}$ state in 12C (Hoyle state), which plays a crucial role in stellar nucleosynthesis Hoyle1954 , lies just above the three-$\alpha$ threshold. It is a resonance with an extremely narrow width of a few electron-volts against decay into three $\alpha$-particles. In such cases, direct NLEFT simulations can reliably extract its energy PRL106-192501 , PRL109-252501 , PRL110-112502 . However, these calculations provide no information on the resonant width and cannot be directly applied to broad resonances. So far, precisely describing general resonance properties remains particularly challenging within NLEFT.
@@
Generally, the single-particle resonances in the potential model correspond to the poles of $S$-matrix. A variety of theoretical approaches have been established to extract resonance parameters, $\it{i.e.}$ resonance energy $E$ and width $\Gamma$. These include the $R$-matrix and $K$-matrix formalisms PR72-29 , PRL59-763 , PRC44-2530 , Taylor2012 , PRL109-072501 , PRC88-024323 , PLB850-138511 based on fitting to scattering or reaction observables, the Green’s function techniques SJNP45-783 , PRC90-054321 , CPC44-084105 , the complex scaling method (CSM) PRC34-95 , PRC37-383 , PRL79-2217 , the complex momentum representation method PRC73-034321 , PRL117-062502 , PLB801-135174 , NST2023Xu , the real stabilization method PRA1-1109 , PRC77-014312 and the phase-shift method PRC2003Sandulescu suitable for very narrow resonance, as well as the practical analytical continuation in the coupling constant (ACCC) method especially suitable for broad resonances PRC70-034308 , EPJA48-40 , PLB730-30 , PRC92-024324 , CCP21-1154 , CPC49-014108 , etc.
[…]
@@
In lattice quantum chromodynamics, Lüscher-type finite-volume methods connect the discrete finite-volume spectrum to scattering amplitudes and resonance parameters, providing a first-principles framework for studying elastic and coupled-channel hadronic resonances PRD83-094505 , PRD87-034505 , PRD91-054008 . Resonances appear as plateaus, with slopes indicating widths. This finite-volume approach was recently applied in NLEFT using Wigner-SU(4) and N3LO chiral forces to study the plausible tetraneutron resonance arXiv:2601.01801 . An efficient alternative involves computing scattering-amplitude poles in the complex-energy plane, where pole positions directly yield resonance energies (real parts) and widths (imaginary parts). These methods provide more direct resonance signatures and better resolve complex structures like overlapping poles or broad resonances. However, they require analytical continuation from real to imaginary momentum, which is a typically ill-posed problem introducing significant numerical uncertainties. Precise calculations thus demand both high-accuracy numerics and efficient analytical-continuation algorithms.